OpenStudy (anonymous):

Solve for x. x^2 + 6x = 0

OpenStudy (anonymous):

\[x^2 + 6x = 0\]\[x(x+6)=0\]can you finish the problem?

OpenStudy (anonymous):

hint: `ZERO PROPERTY `

OpenStudy (anonymous):

what's zero property ?

OpenStudy (anonymous):

x=0,-6

OpenStudy (anonymous):

thanks again (: